OpenStudy (anonymous):

solve for x...but how? (x-7)/5 + 1/5 = -x/10

OpenStudy (phi):

If you add +x/10 to both sides, what do you get?

OpenStudy (anonymous):

everything to the left and it equally 0

OpenStudy (phi):

if we write it out: (x-7)/5 + x/10 +1/5=0 we could do a couple of things. one way: we have fractions. find the common denominator (it is 10) or, another way. If we multiply everything (both sides, all terms) by 10, we can make the denominators go away. which way do you want to go?

OpenStudy (phi):

Both ways are really the same thing, or at least, get us to the same place.

OpenStudy (anonymous):

make the denominators go away

OpenStudy (phi):

so multiply both sides by 10: 10*( (x-7)/5 + x/10 +1/5 ) = 10*0 distribute the 10 (that means multiply 10 times each term inside the parens) what do you get?

OpenStudy (anonymous):

10x-70/50 + 10/50 = -10/-100

OpenStudy (anonymous):

-10x/-100

OpenStudy (phi):

I'll do the first term \[ 10\cdot \frac{(x-7)}{5}= 2(x-7) \] divide the 5 into the 10 to get 2

OpenStudy (phi):

pretty close, but when you multiply by 10 you are multiplying by 10 not 10/10 do it the way I just posted.

OpenStudy (anonymous):

okay

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

10x-70/50 = 2x-14

OpenStudy (phi):

pay attention: we start with 10*( (x-7)/5 + x/10 +1/5 ) = 10*0 don't think too hard. just multiply 10 times each term (that means write 10* next to each term) like this: 10*(x-7)/5 + 10*x/10 + 10*1/5= 10* 0 now divide the bottom into the top can you do that ?

OpenStudy (anonymous):

i really dont get what youre saying

OpenStudy (anonymous):

i dont understand how you got 10 times 0

OpenStudy (phi):

remember we started with (x-7)/5 + x/10 +1/5=0 then we multiply BOTH SIDES by 10 because that is what we have to do to keep things equal. If you multiply 10 times the left side, you do the same to the right (to be fair!) so we write down 10*( (x-7)/5 + x/10 +1/5 ) = 10*0 now on the right side, 10*0 is 0 right? and on the left, it is 10*(x-7)/5 + 10*x/10 + 10*1/5= 0

OpenStudy (anonymous):

yes i understand now

OpenStudy (phi):

now you should learn this 10*(x-7)/5 can be simplified by dividing 5 into 10 to get 2. we can re-write it as 2*(x-7) can you do the other two terms?

OpenStudy (phi):

10*x/10

OpenStudy (anonymous):

x

OpenStudy (anonymous):

2(x-7)+x+2

OpenStudy (phi):

yes, now the last one 10*1/5

OpenStudy (anonymous):

2

OpenStudy (phi):

ok, you went too fast. so you got 2(x-7)+x+2 = 0 (don't leave out the = 0, they are important) can you distribute the 2 (multiply 2 times each term inside the parens)

OpenStudy (anonymous):

2x-14+x+2=0

OpenStudy (phi):

yes. now 2x+1x is ?

OpenStudy (anonymous):

3x

OpenStudy (phi):

and -14+2 is ?

OpenStudy (anonymous):

-12.... i get it now x-4

OpenStudy (phi):

?

OpenStudy (anonymous):

huh....it's 2x+x=12 then 3x=12 x=4

OpenStudy (phi):

yes. That is how you solve it.

OpenStudy (anonymous):

thank you

OpenStudy (phi):

we can check it, by replacing x with 4 in (x-7)/5 + 1/5 = -x/10 (4-7)/5 + 1/5= -4/10 (4-7) is -3, so -3/5 +1/5 = -4/10 -2/5 = -4/10 simplify the right side by dividing the top by 2 and the bottom by 2 -2/5 = -2/5 it worked. Good job.

OpenStudy (anonymous):

yes...thank you!

hero (hero):

(x-7)/5 + 1/5 = -x/10 Given: \(\large\frac{x-7}{5} + \frac{1}{5} = -\frac{x}{10}\) Combine fractions on the left side to get: \(\large\frac{x - 6}{5} = -\frac{x}{10}\) Cross Multiply to get: \(10(x-6) = -5x\) Distribute the 10 across (x-6) to get: 10x - 60 = -5x Add 5x to both sides; add 60 to both sides: 10x + 5x = 60 Simply the left side: 15x = 60 Divide both sides by 15 x = 4